OpenStudy (reneeshumpert):

What is the x-coordinate of the solution of the following system of equations? 3x + y = 6 x - y = 6 -3 0 3 I believe the answer is 0 ?

OpenStudy (nincompoop):

how is that answer zero? SHOW YOUR SOLUTION

OpenStudy (nincompoop):



OpenStudy (texaschic101):

x - y = 6...-y = -x + 6.....y = x - 6 3x + y = 6 3x + x - 6 = 6 4x = 6 + 6 4x = 12 x = 3 x - y = 6 3 - y = 6 -y = 6 - 3 -y = 3 y = -3 check... 3x + y = 6 3(3) - 3 = 6 9 - 3 = 6 6 = 6 (correct) x intercept is (3,0)

OpenStudy (e.mccormick):

The simplest way with this pair of equations is elimination: 3x + y = 6 x - y = 6 Add the second equation to the first: 3x +x + y - y = 6 +6 4x = 12 x=3 Therefore, the x part of the solution is 3 and not 0.